OpenStudy (anonymous):

Find a line that has a y-intercept of 6 and is parallel to the line of 2x+3y+4=0.

OpenStudy (anonymous):

first you have to put the equation in slope intercept form 2x+3y+4=0 3y+4=-2x 3y=-2x-4 y=-2/3x-4/3 notice that the slope is -2/3 having a line parallel means it has the same slope now put it in point slope form: y- ysub 1=m(x-x sub 1) (it has to have a y value of 6) y-6=-2/3(x-0) solve y=-2/3x+6 graph and check to make sure it is parallel

OpenStudy (anonymous):

ook thanks!